OpenStudy (anonymous):

2x^2−(2√2+1)x+√2=0 solve using the quadratic formula

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

lol this equation again huh? :)

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

i started it but i can't simplify it.

zepdrix (zepdrix):

\[\Large\rm x=\frac{-\color{orangered}{b}\pm\sqrt{\color{orangered}{b}^2-4\color{royalblue}{a}\color{#CC0033}{c}}}{2\color{royalblue}{a}}\] And we have:\[\Large\rm \color{royalblue}{2}x^2+\color{orangered}{-(2\sqrt2+1)}x+\color{#CC0033}{\sqrt2}=0\]

OpenStudy (anonymous):

yep i did plug in all the numbers into the formula

OpenStudy (anonymous):

i just get so many radicals in the sqaureroot

zepdrix (zepdrix):

\[\Large\rm x=\frac{\color{orangered}{2\sqrt2+1}\pm\sqrt{\color{orangered}{(2\sqrt2+1)}^2-4\color{royalblue}{(2)}\color{#CC0033}{(\sqrt2)}}}{2\color{royalblue}{(2)}}\]Having trouble simplifying it or something? :o

zepdrix (zepdrix):

Hmmm

zepdrix (zepdrix):

\[\Large\rm x=\frac{2\sqrt2+1\pm\sqrt{8+4\sqrt2+1-8\sqrt2}}{4}\]Oooo I thought those sqrt(2) would cancel out under the root :( Hmmm

OpenStudy (anonymous):

exactly thats where i got stuck

zepdrix (zepdrix):

\[\Large\rm x=\frac{2\sqrt2+1\pm\sqrt{9-4\sqrt2}}{4}\]

zepdrix (zepdrix):

Wolfram is telling me that \(\Large\rm \sqrt{9-4\sqrt2}\) is equivalent to \(\Large\rm 2\sqrt{2}-1\). Hmm I'm trying to figure out how we get there...

OpenStudy (anonymous):

oh

zepdrix (zepdrix):

Hmm there is a method that will work for `Denesting` square roots. But it's a little weird :o Why didn't you use the grouping method? No bueno? That works out much much easier.

OpenStudy (anonymous):

well we're required to use grouping, quadratic formula, and completing the square, and graphing

zepdrix (zepdrix):

Ok I'm reading a website that explains denesting square roots. I'll post the method, it doesn't seem too bad.

zepdrix (zepdrix):

Try not to confuse these letters with our quadratic formula letters :) So we have something of this form: \[\Large\rm \sqrt{9-4\sqrt2}=\sqrt{a-b\sqrt{c}}\] We first calculate: \(\Large\rm a^2-b^2\cdot c\) If this value turns out to be a perfect square, then we are able to simplify our expression. So for our problem we get: \(\Large\rm 9^2-4^2\cdot 2=49\) Which is \(\Large\rm 7^2\) and we'll call this \(\Large\rm q^2\). So \(\Large\rm 7=q\). So we're allowed to rewrite our expression like this: \[\Large\rm \sqrt{a-b\sqrt{c}}=\sqrt{\frac{a+q}{2}}+\sqrt{\frac{a-q}{2}}\]So for our problem:\[\Large\rm \sqrt{9-4\sqrt{2}} =\sqrt{\frac{9+7}{2}}+\sqrt{\frac{9-7}{2}}\]

zepdrix (zepdrix):

Which simplifies a bit further, yes? :o \[\Large\rm =\sqrt{8}+\sqrt{1}\]\[\Large\rm =2\sqrt{2}+1\]

OpenStudy (anonymous):

but how do you get 1/2 and square root 2

zepdrix (zepdrix):

So that only takes care of this orange part:\[\Large\rm x=\frac{2\sqrt2+1\pm\color{orangered}{\sqrt{9-4\sqrt2}}}{4}\] Which becomes:\[\Large\rm x=\frac{2\sqrt2+1\pm\color{orangered}{(2\sqrt2+1)}}{4}\]

zepdrix (zepdrix):

And now you need to look at the plus/minus separately to get your two solutions.\[\Large\rm x=\frac{2\sqrt2+1\color{royalblue}{\pm}(2\sqrt2+1)}{4}\]Turns into two equations:\[\Large\rm x=\frac{2\sqrt2+1\color{royalblue}{+}(2\sqrt2+1)}{4}\]and\[\Large\rm x=\frac{2\sqrt2+1\color{royalblue}{-}(2\sqrt2+1)}{4}\]

zepdrix (zepdrix):

AHhhh I did the wrong simplification for the root :( It can also be written as: \(\Large\rm =\sqrt{\frac{a+q}{2}}-\sqrt{\frac{a-q}{2}}\) That's the one we wanted ^ With the minus between the terms. Lemme fix that in our setup.

zepdrix (zepdrix):

Ok so these are our two equations: \[\Large\rm x=\frac{2\sqrt2+1\color{royalblue}{+}(2\sqrt2-1)}{4}\]and\[\Large\rm x=\frac{2\sqrt2+1\color{royalblue}{-}(2\sqrt2-1)}{4}\]

zepdrix (zepdrix):

Do you see how the `1`'s will cancel out in the `plus` equation, and how the `2sqrt(2)`'s will cancel out in the `minus` equation?

zepdrix (zepdrix):

Quadratic Formula REALLY sucks for these types of roots :O( lol

OpenStudy (anonymous):

yelp could you also help solve it completing the square and graping(wolfram)/

zepdrix (zepdrix):

What type of class is this for? If it's for a Calculus class, we can get a more accurate graph by finding the max/min values. Otherwise if it's only for Algebra or something, we can graph the turning points and then use a graphing calculator or something to rough in the rest.

zepdrix (zepdrix):

Completing the square? Oh boy, that's going to be awful lolol

zepdrix (zepdrix):

\[\Large\rm 2x^2-(2\sqrt2+1)x+\sqrt2=0\]Subtract sqrt2 to the other side, then divide both sides by 2,\[\Large\rm x^2-\left(\frac{2\sqrt2+1}{2}\right)x\qquad\qquad\qquad=-\frac{\sqrt{2}}{2}\]To complete the square, we take half of our b coefficient, and square it. That's the value we want to add to each side.\[\Large\rm \left(\frac{b}{2}\right)^2=\left(-\frac{2\sqrt2+1}{4}\right)^2=\left(\frac{2\sqrt2+1}{4}\right)^2\]So we add that to each side,\[\large\rm x^2-\left(\frac{2\sqrt2+1}{2}\right)x+\left(\frac{2\sqrt2+1}{4}\right)^2=-\frac{\sqrt{2}}{2}+\left(\frac{2\sqrt2+1}{4}\right)^2\]Our perfect square will condense down to \(\Large\rm \left(x+\frac{b}{2}\right)^2\), giving us,\[\large\rm \left(x-\frac{2\sqrt2+1}{4}\right)^2=-\frac{\sqrt{2}}{2}+\left(\frac{2\sqrt2+1}{4}\right)^2\]

zepdrix (zepdrix):

And then do some square root business,\[\large\rm x-\frac{2\sqrt2+1}{4}=\pm\sqrt{-\frac{\sqrt{2}}{2}+\left(\frac{2\sqrt2+1}{4}\right)^2}\]

zepdrix (zepdrix):

And then some more steps and some things.... It'll probably work out similar to the Quadratic Formula method from here. I dunno, I'm too tired :d Time to call it a night I think >.<